Ask
your own question, for FREE!
Mathematics
17 Online
You draw two cards at random from a standard deck. What is the probability of drawing at least one diamond?
Still Need Help?
Join the QuestionCove community and study together with friends!
A. 9/34 B. 3/4 C. 9/16 D. 15/34
thank god I"ve been stuck on this question for 2 hours its very hard for me
so 12/54 right?
nopes how u got that?
P(atleast 1 diamond) = 1- P(No diamond)
Still Need Help?
Join the QuestionCove community and study together with friends!
someone joined earlier and sais that
P (diamond) = 3/4
ok
u get how P(diamond) = 3/4?
so u think c
Still Need Help?
Join the QuestionCove community and study together with friends!
one guy said d
but i do get it more then normally
i dont tihnk its 3/4
why?
it was wrong
Still Need Help?
Join the QuestionCove community and study together with friends!
but thxs for actually helping me
i never said 3/4 is correct i just asked you why you think its not correct? :P
P(No Diamond) = (3/4)*(38/51) = 19/34
P(atleast 1 diamond) = 1-19/34 =...
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
15 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal